\begin{thebibliography}{99}



\bibitem{Tsallis1988}
C.~Tsallis, ``{Possible generalization of Boltzmann-Gibbs statistics.}'',
  \href{https://doi.org/10.1007/BF01016429}{\emph{J. Stat. Phys.} {\bfseries 52}
  (1988) 479--487}.

\bibitem{tsallis2003}
C.~Tsallis, F.~Baldovin, R.~Cerbino and P.~Pierobon, ``Introduction to
  nonextensive statistical mechanics and thermodynamics'',
  \href{https://arxiv.org/abs/cond-mat/0309093}{{\ttfamily
  arXiv:cond-mat/0309093}}.

\bibitem{ABULMAGD2007450}
A.~Abul-Magd, ``Nonextensive random-matrix theory based on kaniadakis
  entropy'',
  \href{https://doi.org/https://doi.org/10.1016/j.physleta.2006.09.080}{\emph{Phys. Lett. A} {\bfseries 361} (2007) 450--454}.

\bibitem{e17010052}
R.~V. Chamberlin, ``{The Big World of Nanothermodynamics}'',
  \href{https://doi.org/10.3390/e17010052}{\emph{Entropy} {\bfseries 17} (2015)
  52--73}, [\href{https://arxiv.org/abs/1504.04754}{{\ttfamily
  arXiv:1504.04754}}].

\bibitem{Nivanen_2003Kaniadakis}
L.~Nivanen, A.~{Le Méhauté} and Q.~Wang, ``Generalized algebra within a
  nonextensive statistics'',
  \href{https://doi.org/https://doi.org/10.1016/S0034-4877(03)80040-X}{\emph{Rep. Math. Phys.} {\bfseries 52} (2003) 437--444}.

\bibitem{Chamblin:1999tk}
A.~Chamblin, R.~Emparan, C.~V. Johnson and R.~C. Myers, ``{Charged AdS black
  holes and catastrophic holography}'',
  \href{https://doi.org/10.1103/PhysRevD.60.064018}{\emph{Phys. Rev. D}
  {\bfseries 60} (1999) 064018},
  [\href{https://arxiv.org/abs/hep-th/9902170}{{\ttfamily
  arXiv:hep-th/9902170}}].

\bibitem{Chamblin:1999hg}
A.~Chamblin, R.~Emparan, C.~V. Johnson and R.~C. Myers, ``{Holography,
  thermodynamics and fluctuations of charged AdS black holes}'',
  \href{https://doi.org/10.1103/PhysRevD.60.104026}{\emph{Phys. Rev. D}
  {\bfseries 60} (1999) 104026},
  [\href{https://arxiv.org/abs/hep-th/9904197}{{\ttfamily
  arXiv:hep-th/9904197}}].

\bibitem{Kubiznak:2012wp}
D.~Kubiznak and R.~B. Mann, ``{P-V criticality of charged AdS black holes}'',
  \href{https://doi.org/10.1007/JHEP07(2012)033}{\emph{JHEP} {\bfseries 07}
  (2012) 033}, [\href{https://arxiv.org/abs/1205.0559}{{\ttfamily
  arXiv:1205.0559}}].

\bibitem{Kastor:2009wy}
D.~Kastor, S.~Ray and J.~Traschen, ``{Enthalpy and the Mechanics of AdS Black
  Holes}'', \href{https://doi.org/10.1088/0264-9381/26/19/195011}{\emph{Class.
  Quant. Grav.} {\bfseries 26} (2009) 195011},
  [\href{https://arxiv.org/abs/0904.2765}{{\ttfamily arXiv:0904.2765}}].

\bibitem{Dolan:2010ha}
B.~P. Dolan, ``{The cosmological constant and the black hole equation of
  state}'', \href{https://doi.org/10.1088/0264-9381/28/12/125020}{\emph{Class.
  Quant. Grav.} {\bfseries 28} (2011) 125020},
  [\href{https://arxiv.org/abs/1008.5023}{{\ttfamily arXiv:1008.5023}}].

\bibitem{Dolan:2011xt}
B.~P. Dolan, ``{Pressure and volume in the first law of black hole
  thermodynamics}'',
  \href{https://doi.org/10.1088/0264-9381/28/23/235017}{\emph{Class. Quant.
  Grav.} {\bfseries 28} (2011) 235017},
  [\href{https://arxiv.org/abs/1106.6260}{{\ttfamily arXiv:1106.6260}}].

\bibitem{Cvetic:2010jb}
M.~Cvetic, G.~W. Gibbons, D.~Kubiznak and C.~N. Pope, ``{Black Hole Enthalpy
  and an Entropy Inequality for the Thermodynamic Volume}'',
  \href{https://doi.org/10.1103/PhysRevD.84.024037}{\emph{Phys. Rev. D}
  {\bfseries 84} (2011) 024037},
  [\href{https://arxiv.org/abs/1012.2888}{{\ttfamily arXiv:1012.2888}}].

\bibitem{Frassino:2015oca}
A.~M. Frassino, R.~B. Mann and J.~R. Mureika, ``{Lower-Dimensional Black Hole
  Chemistry}'', \href{https://doi.org/10.1103/PhysRevD.92.124069}{\emph{Phys.
  Rev. D} {\bfseries 92} (2015) 124069},
  [\href{https://arxiv.org/abs/1509.05481}{{\ttfamily arXiv:1509.05481}}].

\bibitem{Kubiznak:2016qmn}
D.~Kubiznak, R.~B. Mann and M.~Teo, ``{Black hole chemistry: thermodynamics
  with Lambda}'', \href{https://doi.org/10.1088/1361-6382/aa5c69}{\emph{Class.
  Quant. Grav.} {\bfseries 34} (2017) 063001},
  [\href{https://arxiv.org/abs/1608.06147}{{\ttfamily arXiv:1608.06147}}].

\bibitem{Mann:2025xrb}
R.~B. Mann, ``{Black hole chemistry: The first 15 years}'',
  \href{https://doi.org/10.1142/S0218271825420015}{\emph{Int. J. Mod. Phys. D}
  {\bfseries 34} (2025) 2542001}.

\bibitem{Strominger:1996sh}
A.~Strominger and C.~Vafa, ``{Microscopic origin of the Bekenstein-Hawking
  entropy}'', \href{https://doi.org/10.1016/0370-2693(96)00345-0}{\emph{Phys.
  Lett. B} {\bfseries 379} (1996) 99--104},
  [\href{https://arxiv.org/abs/hep-th/9601029}{{\ttfamily
  arXiv:hep-th/9601029}}].

\bibitem{Maldacena:1996gb}
J.~M. Maldacena and A.~Strominger, ``{Statistical entropy of four-dimensional
  extremal black holes}'',
  \href{https://doi.org/10.1103/PhysRevLett.77.428}{\emph{Phys. Rev. Lett.}
  {\bfseries 77} (1996) 428--429},
  [\href{https://arxiv.org/abs/hep-th/9603060}{{\ttfamily
  arXiv:hep-th/9603060}}].

\bibitem{Wei:2015iwa}
S.-W. Wei and Y.-X. Liu, ``{Insight into the Microscopic Structure of an AdS
  Black Hole from a Thermodynamical Phase Transition}'',
  \href{https://doi.org/10.1103/PhysRevLett.115.111302}{\emph{Phys. Rev. Lett.}
  {\bfseries 115} (2015) 111302},
  [\href{https://arxiv.org/abs/1502.00386}{{\ttfamily arXiv:1502.00386}}].

\bibitem{Kastor:2014dra}
D.~Kastor, S.~Ray and J.~Traschen, ``{Chemical Potential in the First Law for
  Holographic Entanglement Entropy}'',
  \href{https://doi.org/10.1007/JHEP11(2014)120}{\emph{JHEP} {\bfseries 11}
  (2014) 120}, [\href{https://arxiv.org/abs/1409.3521}{{\ttfamily
  arXiv:1409.3521}}].

\bibitem{Zhang:2014uoa}
J.-L. Zhang, R.-G. Cai and H.~Yu, ``{Phase transition and thermodynamical
  geometry for Schwarzschild AdS black hole in AdS$_{5}$ {\texttimes} S$^{5}$
  spacetime}'', \href{https://doi.org/10.1007/JHEP02(2015)143}{\emph{JHEP}
  {\bfseries 02} (2015) 143},
  [\href{https://arxiv.org/abs/1409.5305}{{\ttfamily arXiv:1409.5305}}].

\bibitem{Karch:2015rpa}
A.~Karch and B.~Robinson, ``{Holographic Black Hole Chemistry}'',
  \href{https://doi.org/10.1007/JHEP12(2015)073}{\emph{JHEP} {\bfseries 12}
  (2015) 073}, [\href{https://arxiv.org/abs/1510.02472}{{\ttfamily
  arXiv:1510.02472}}].

\bibitem{Visser:2021eqk}
M.~R. Visser, ``{Holographic thermodynamics requires a chemical potential for
  color}'', \href{https://doi.org/10.1103/PhysRevD.105.106014}{\emph{Phys. Rev.
  D} {\bfseries 105} (2022) 106014},
  [\href{https://arxiv.org/abs/2101.04145}{{\ttfamily arXiv:2101.04145}}].

\bibitem{Cong:2021fnf}
W.~Cong, D.~Kubiznak and R.~B. Mann, ``{Thermodynamics of AdS Black Holes:
  Critical Behavior of the Central Charge}'',
  \href{https://doi.org/10.1103/PhysRevLett.127.091301}{\emph{Phys. Rev. Lett.}
  {\bfseries 127} (2021) 091301},
  [\href{https://arxiv.org/abs/2105.02223}{{\ttfamily arXiv:2105.02223}}].

\bibitem{Cong:2021jgb}
W.~Cong, D.~Kubiznak, R.~B. Mann and M.~R. Visser, ``{Holographic CFT phase
  transitions and criticality for charged AdS black holes}'',
  \href{https://doi.org/10.1007/JHEP08(2022)174}{\emph{JHEP} {\bfseries 08}
  (2022) 174}, [\href{https://arxiv.org/abs/2112.14848}{{\ttfamily
  arXiv:2112.14848}}].

\bibitem{Ahmed:2023dnh}
M.~B. Ahmed, W.~Cong, D.~Kubiznak, R.~B. Mann and M.~R. Visser, ``{Holographic
  CFT phase transitions and criticality for rotating AdS black holes}'',
  \href{https://doi.org/10.1007/JHEP08(2023)142}{\emph{JHEP} {\bfseries 08}
  (2023) 142}, [\href{https://arxiv.org/abs/2305.03161}{{\ttfamily
  arXiv:2305.03161}}].

\bibitem{Ahmed:2023snm}
M.~B. Ahmed, W.~Cong, D.~Kubiz{\v{n}}{\'a}k, R.~B. Mann and M.~R. Visser,
  ``{Holographic Dual of Extended Black Hole Thermodynamics}'',
  \href{https://doi.org/10.1103/PhysRevLett.130.181401}{\emph{Phys. Rev. Lett.}
  {\bfseries 130} (2023) 181401},
  [\href{https://arxiv.org/abs/2302.08163}{{\ttfamily arXiv:2302.08163}}].

\bibitem{Punia:2023ilo}
G.~S.~Punia,
``Bulk-boundary thermodynamics of charged black holes in higher-derivative theory,''
    \href{https://doi.org/10.1140/epjc/s10052-024-12753-2}{\emph{Eur. Phys. J. C}
  {\bfseries 84} (2024) 467},
  [\href{https://arxiv.org/abs/2305.06552}{{\ttfamily arXiv:2305.06552}}].

\bibitem{Yang:2024krx}
S.-J. Yang, M.~S. Ali, S.-W. Wei and Y.-X. Liu, ``{Holographic thermodynamics
  of a five-dimensional neutral Gauss{\textendash}Bonnet AdS black hole}'',
  \href{https://doi.org/10.1140/epjc/s10052-025-14136-7}{\emph{Eur. Phys. J. C}
  {\bfseries 85} (2025) 420},
  [\href{https://arxiv.org/abs/2408.00318}{{\ttfamily arXiv:2408.00318}}].

\bibitem{Ladghami:2024wkv}
Y.~Ladghami and T.~Ouali, ``{Black holes thermodynamics with CFT re-scaling}'',
  \href{https://doi.org/10.1016/j.dark.2024.101471}{\emph{Phys. Dark Univ.}
  {\bfseries 44} (2024) 101471},
  [\href{https://arxiv.org/abs/2402.15913}{{\ttfamily arXiv:2402.15913}}].

\bibitem{Zhang:2025dgm}
M.~Zhang, W.-D. Tan, M.~Lu, D.~Bhattacharya, J.~Yang and R.~B. Mann,
  ``{Finite-cutoff Holographic Thermodynamics}'',
  \href{https://doi.org/https://doi.org/10.48550/arXiv.2507.01010} [\href{https://arxiv.org/abs/2507.01010}{{\ttfamily
  arXiv:2507.01010}}].

\bibitem{Zeyuan:2021uol}
G.~Zeyuan and L.~Zhao, ``{Restricted phase space thermodynamics for AdS black
  holes via holography}'',
  \href{https://doi.org/10.1088/1361-6382/ac566c}{\emph{Class. Quant. Grav.}
  {\bfseries 39} (2022) 075019},
  [\href{https://arxiv.org/abs/2112.02386}{{\ttfamily arXiv:2112.02386}}].

\bibitem{Zhao:2022dgc}
L.~Zhao, ``{Thermodynamics for higher dimensional rotating black holes with
  variable Newton constant *}'',
  \href{https://doi.org/10.1088/1674-1137/ac4f4c}{\emph{Chin. Phys. C}
  {\bfseries 46} (2022) 055105},
  [\href{https://arxiv.org/abs/2201.00521}{{\ttfamily arXiv:2201.00521}}].

\bibitem{Gao:2021xtt}
Z.~Gao, X.~Kong and L.~Zhao, ``{Thermodynamics of Kerr-AdS black holes in the
  restricted phase space}'',
  \href{https://doi.org/10.1140/epjc/s10052-022-10080-y}{\emph{Eur. Phys. J. C}
  {\bfseries 82} (2022) 112},
  [\href{https://arxiv.org/abs/2112.08672}{{\ttfamily arXiv:2112.08672}}].

\bibitem{Wald:1993nt}
R.~M. Wald, ``{Black hole entropy is the Noether charge}'',
  \href{https://doi.org/10.1103/PhysRevD.48.R3427}{\emph{Phys. Rev. D}
  {\bfseries 48} (1993) R3427--R3431},
  [\href{https://arxiv.org/abs/gr-qc/9307038}{{\ttfamily
  arXiv:gr-qc/9307038}}].

\bibitem{Iyer:1994ys}
V.~Iyer and R.~M. Wald, ``{Some properties of Noether charge and a proposal for
  dynamical black hole entropy}'',
  \href{https://doi.org/10.1103/PhysRevD.50.846}{\emph{Phys. Rev. D} {\bfseries
  50} (1994) 846--864}, [\href{https://arxiv.org/abs/gr-qc/9403028}{{\ttfamily
  arXiv:gr-qc/9403028}}].

\bibitem{Xiao:2023lap}
Y.~Xiao, Y.~Tian and Y.-X. Liu, ``{Extended Black Hole Thermodynamics from
  Extended Iyer-Wald Formalism}'',
  \href{https://doi.org/10.1103/PhysRevLett.132.021401}{\emph{Phys. Rev. Lett.}
  {\bfseries 132} (2024) 021401},
  [\href{https://arxiv.org/abs/2308.12630}{{\ttfamily arXiv:2308.12630}}].

\bibitem{Hajian:2023bhq}
K.~Hajian and B.~Tekin, ``{Coupling Constants as Conserved Charges in Black
  Hole Thermodynamics}'',
  \href{https://doi.org/10.1103/PhysRevLett.132.191401}{\emph{Phys. Rev. Lett.}
  {\bfseries 132} (2024) 191401},
  [\href{https://arxiv.org/abs/2309.07634}{{\ttfamily arXiv:2309.07634}}].

\bibitem{Cai:2024tyv}
R.-G. Cai, L.~Li and J.-K. Zhao, ``{Thermodynamics of dyonic black holes in
  minimal supergravity}'',
  \href{https://doi.org/https://doi.org/10.48550/arXiv.2410.00717}
  [\href{https://arxiv.org/abs/2410.00717}{{\ttfamily
  arXiv:2410.00717}}].

\bibitem{Campos:2025nnk}
T.~d.~L. Campos, M.~C. Baldiotti and C.~Molina, ``{Black-Hole Thermodynamics
  from Gauge Freedom in Extended Iyer-Wald Formalism}'',
  \href{https://doi.org/10.3390/universe11070215}{\emph{Universe} {\bfseries
  11} (2025) 215}, [\href{https://arxiv.org/abs/2507.03751}{{\ttfamily
  arXiv:2507.03751}}].

\bibitem{Tavlayan:2024zbl}
A.~Tavlayan and B.~Tekin, ``{Refined thermodynamics of black holes with proper
  conserved charges}'',
  \href{https://doi.org/10.1103/PhysRevD.110.084049}{\emph{Phys. Rev. D}
  {\bfseries 110} (2024) 084049},
  [\href{https://arxiv.org/abs/2407.04002}{{\ttfamily arXiv:2407.04002}}].

\bibitem{Wu:2025jes}
S.-P. Wu, Y.-X. Liu and S.-W. Wei, ``{Generalized Free Energy Landscapes from
  Iyer-Wald Formalism}'',
  \href{https://doi.org/https://doi.org/10.48550/arXiv.2504.00503}
  [\href{https://arxiv.org/abs/2504.00503}{{\ttfamily
  arXiv:2504.00503}}].

\bibitem{Lee:1990nz}
J.~Lee and R.~M. Wald, ``{Local symmetries and constraints}'',
  \href{https://doi.org/10.1063/1.528801}{\emph{J. Math. Phys.} {\bfseries 31}
  (1990) 725--743}.

\bibitem{Myers:2010tj}
R.~C. Myers and A.~Sinha, ``{Holographic c-theorems in arbitrary dimensions}'',
  \href{https://doi.org/10.1007/JHEP01(2011)125}{\emph{JHEP} {\bfseries 01}
  (2011) 125}, [\href{https://arxiv.org/abs/1011.5819}{{\ttfamily
  arXiv:1011.5819}}].

\bibitem{york1986black}
J.~W. York~Jr, ``Black-hole thermodynamics and the euclidean einstein action'',
  \href{https://doi.org/10.1103/PhysRevD.33.2092}{\emph{Phys. Rev. D}
  {\bfseries 33} (1986) 2092}.

\bibitem{Hawking:1995ap}
S.~W.~Hawking and S.~F.~Ross,
``Duality between electric and magnetic black holes,''
\href{https://doi.org/10.1103/PhysRevD.52.5865}{\emph{Phys. Rev. D} {\bfseries 52}
  (1995) 5865}, [\href{https://arxiv.org/abs/hep-th/9504019}{{\ttfamily
  arXiv:hep-th/9504019}}].

\bibitem{Gibbons:2004ai}
G.~W. Gibbons, M.~J. Perry and C.~N. Pope, ``{The First law of thermodynamics
  for Kerr-anti-de Sitter black holes}'',
  \href{https://doi.org/10.1088/0264-9381/22/9/002}{\emph{Class. Quant. Grav.}
  {\bfseries 22} (2005) 1503--1526},
  [\href{https://arxiv.org/abs/hep-th/0408217}{{\ttfamily
  arXiv:hep-th/0408217}}].
  
  
\bibitem{Li:2018drw}
Y.~Z.~Li, H.~L{\"u} and Z.~F.~Mai,
``Universal Structure of Covariant Holographic Two-Point Functions In Massless Higher-Order Gravities,''
\href{https://doi.org/10.1007/JHEP10(2018)063}{\emph{JHEP} {\bfseries 10} 
  (2018) 063},
  [\href{https://arxiv.org/abs/hep-th/1808.00494}{{\ttfamily
  arXiv:1808.00494}}].
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsfonts}
\usepackage{bbm}
\usepackage{graphicx}
\usepackage{pgfplots}
\usepackage{booktabs}
\usepackage{subfigure}
\usepackage{mathrsfs}
\usepackage{color}
\usepackage{verbatim}
\usepackage{cancel}
\usepackage{geometry}             
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{epstopdf}
\usepackage{cases}
\usepackage{cite}
\usepackage{setspace}

\usepackage[hyperindex=true,
pdfstartview=FitH,
bookmarksnumbered=true,
bookmarksopen=true,
citecolor=blue,
linkcolor=blue,
colorlinks=true,
urlcolor=rossoCP3,
unicode]{hyperref}

\definecolor{rossoCP3}{cmyk}{0,.88,.77,.40}

\begin{document}

\title{\bf
Black hole thermodynamics from Iyer-Wald formalism with variable Newton's constant}
{\author{\small Xuan-Rui Chen${}^{1}$, Bin Wu${}^{1,2,3}$\thanks{{\em email}: \href{mailto:binwu@nwu.edu.cn}{binwu@nwu.edu.cn}}{ }, Zhen-Ming Xu${}^{1,2,3}$
        \vspace{5pt}\\
        \small $^{1}${\it School of Physics, Northwest University, Xi'an 710127, China}\\
        \small $^{2}${\it Shaanxi Key Laboratory for Theoretical Physics Frontiers, Xi'an 710127, China}\\
        \small $^{3}${\it Peng Huanwu Center for Fundamental Theory, Xi'an 710127, China}}
	
\date{}
\maketitle
\begin{spacing}{1.15}
\begin{abstract}
Recent advances in black hole thermodynamics have intensified efforts to investigate its thermodynamic framework analogues to that of classical thermal systems. Departing from the extended phase space approach which involves a variable cosmological constant, the restricted phase space formalism focusing on the variations of Newton's gravitational constant in black hole thermodynamics is proposed. This approach enables the construction of a self-consistent thermodynamic structure that features both the first law and the Euler relation. In this paper, we demonstrate that for any diffeomorphism-invariant gravitational theory, the Iyer-Wald formalism is applicable to restricted phase space thermodynamics and construct a self-consistent extensive black hole thermodynamic system. Our work thereby establishes the Iyer-Wald approach as the geometric foundation for restricted phase space black hole thermodynamics.
\end{abstract}

\section{Introduction}
As a universal theory, thermodynamics is widely employed in various condensed matter systems, celestial bodies, and even quantum systems. Especially since the establishment and improvement of the Gibbs ensemble theory, people have successfully linked the macroscopic thermodynamics phenomena to the microscopic mechanism of the system. In equilibrium thermodynamics (or classical thermodynamics), the extensive properties of the system are one of its core assumptions. However, when the system scale is small, the interface effect becomes significant, and for long-range interactions system, it is necessary to move beyond the Gibbs ensemble theory framework to discuss this kind of non-extensive system using Tsallis statistics\cite{Tsallis1988,tsallis2003} or other generalized statistics\cite{ABULMAGD2007450, e17010052,Nivanen_2003Kaniadakis}.

Based on the universality of thermodynamics, it is natural to apply it to black hole systems to analyze their thermal properties. Intriguingly, it is found that the AdS charged black hole exhibits a phase transition analogous to the liquid-gas phase transition in the Van der Waals system\cite{Chamblin:1999tk,Chamblin:1999hg,Kubiznak:2012wp}. Furthermore, when the negative cosmological constant $\Lambda$ is interpreted as thermal pressure, it has been discovered that black holes exhibit rich thermodynamic behaviors\cite{Kastor:2009wy,Dolan:2010ha,Dolan:2011xt,Cvetic:2010jb}. This framework is called extended phase space black hole thermodynamics, and is now also known as black hole chemistry \cite{Frassino:2015oca,Kubiznak:2016qmn}. The literature \cite{Mann:2025xrb} provides a good overview of the black hole chemistry, and the latest developments can be referenced.

However, due to the unknown microscopic components of black holes and the complexity of the internal structure, it is still difficult to identify the microscopic statistical foundation behind the thermal properties of black holes. At present, the research status of black hole thermodynamics is somewhat similar to the era of Boltzmann's research on thermodynamics, during which the existence of atoms and molecules has not been directly confirmed by experiments. In the literature \cite{Strominger:1996sh,Maldacena:1996gb,Wei:2015iwa}, many efforts have been made to study the microscopic statistical of the black hole thermodynamics, obtaining many meaningful results, such as Bekenstein entropy. 

As important as the issue of the microscopic statistical mechanism of black holes, there is another significant topic of whether black hole thermodynamics is consistent with classical thermodynamics or Gibbs ensemble theory. With the development of the AdS/CFT correspondence, it is conjectured that the AdS black hole thermodynamics properties can be understood from a perspective of holography \cite{Kastor:2014dra,Zhang:2014uoa,Karch:2015rpa,Visser:2021eqk,Cong:2021fnf,Cong:2021jgb,Ahmed:2023dnh,Ahmed:2023snm,Punia:2023ilo,Yang:2024krx,Ladghami:2024wkv,Zhang:2025dgm}. The Euler-like relation between the thermal quantities of the dual CFT, which is a central relation in classical thermodynamics, is restored. Since the central charge of the conformal field is related to the cosmological constant and Newton's gravitational constant, restricted phase space thermodynamics emerges by fixing $\Lambda$ while varying Newton's gravitational constant $G$\cite{Zeyuan:2021uol,Zhao:2022dgc,Gao:2021xtt}. The first law and Euler relation of the extensive thermodynamics established within this framework are consistent with each other. 

Moreover, it is worth noting that the Iyer-Wald formalism has proven invaluable in verifying the fundamental relations in black hole thermodynamics\cite{Wald:1993nt,Iyer:1994ys}. This formalism provides a Lagrangian based framework for constructing Noether currents and charges related to spacetime symmetries. Extensions of Iyer-Wald formalism were employed successfully to numerous scenarios, including the extended phase space thermodynamics\cite{Xiao:2023lap}, gravity with couplings of spacetime fields\cite{Hajian:2023bhq}, dyonic black holes in Einstein-Maxwell-Chern-Simons gravity \cite{Cai:2024tyv}, exact isohomogeneous transformation-based\cite{Campos:2025nnk}, maximally symmetric vacuum solution \cite{Tavlayan:2024zbl}, and generalized free energy landscapes \cite{Wu:2025jes}.

Despite these advances, a fundamental geometric foundation for the restricted phase space, which provides important insights into the extensive nature of black hole thermodynamics, remains elusive.
Current approaches often rely on thermodynamic relations $\mu\equiv\left(\partial{M}/\partial{C}\right)_{S,J...}$ or are derived indirectly from the Euclidean action, lacking a first-principles derivation based on diffeomorphism invariance. In this paper, we attempt to employ an extension of the Iyer-Wald formalism to establish the foundational geometry for restricted phase space thermodynamics.

The organization of this paper is as follows: In Section \ref{II}, we briefly review the Iyer-Wald formalism, which effectively establishes the thermodynamic relationships for black holes. Furthermore, we analyze the thermodynamic properties in the phase space where the Newton's constant $G$ varies based on an Iyer-Wald formalism extension. Then, taking the Schwarzschild-AdS black hole, Reissner-Nordström-AdS black hole, and Kerr-AdS black hole models as examples, we derive the homogeneous properties of black hole thermodynamics. Finally, our conclusions are presented in Section \ref{IV}.

\section{Iyer-Wald formalism}\label{II}
The Iyer-Wald formalism is well established, that for any diffeomorphism-invariant gravitational theory, the laws of black hole thermodynamics can be readily derived. In the following, we briefly review the key results, and for further details, see \cite{Wald:1993nt,Iyer:1994ys} and the references therein.

Following the standard procedure of the Iyer-Wald formalism, the variation of the Lagrangian  $\mathbf{L}$ with respect to the dynamical fields $\phi \equiv \{g_{\mu \nu}, \psi\}$ can be written as
\begin{align}
	\delta \mathbf{L}=\mathbf{E}^{\phi}\delta  \phi+ d \boldsymbol{\Theta}[\phi, \delta \phi], \label{1}
\end{align}
where $\mathbf{E}^{\phi}$ represents the equations of motion derived from the Lagrangian, and
$\boldsymbol{\Theta}(\phi, \delta \phi)$ is a total derivative term. A Noether current $\mathbf{J}_\xi$ for an arbitrary vector $\xi$ is defined as follows
\begin{align}
	\mathbf{J}_\xi \equiv \boldsymbol{\Theta}[\phi, \mathcal{L}_\xi \phi]-\xi  \cdot \mathbf{L},
	\label{NC}
\end{align}
which satisfies $d\mathbf{J}_\xi=-\mathbf{E}^\phi \mathcal{L}_\xi \phi$\cite{Lee:1990nz}. Thus, a Noether charge can be constructed from $\mathbf{J}_\xi= d \mathbf{Q}_\xi$ under the on-shell condition.

By combining the variation of Equation (\ref{NC}) with the Killing symmetry requirement for $\xi$ and integrating the resulting expression over a hypersurface $\Sigma$, we derive the foundational thermodynamic relations governing black hole thermodynamics.
\begin{align}
	\int_{\partial \Sigma} &\left(  \delta \mathbf{Q}_{\xi} -  \xi \cdot \boldsymbol{\Theta}(\phi, \delta \phi) \right)  =0, \label{intfor1} \\
	& \int_{\partial \Sigma} \mathbf{Q}_{\xi} = -\int_{\Sigma}\xi  \cdot \mathbf{L}. \label{intfor2}
\end{align}
The horizon $S_H$ functions as an intrinsic integration boundary where the Killing vector field $\xi$ vanishes, namely $\xi_H=0$. The other integration boundary may take the form of any surface $S_r$ with an arbitrary radius surrounding the black hole. Later, we will notice that the specific radial value $r$ of the surface, irrespective of its location, has no bearing on the results. In particular, within the framework of Wald entropy, it can be shown that the relations $ \int_{S_H} \delta \mathbf{Q}_{\xi_H}=T\delta S$ and $\int_{S_H} \mathbf{Q}_{\xi_H} = TS$ hold \cite{Wald:1993nt,Iyer:1994ys}.

In this paper, our primary goal is to provide the formulation of the restricted phase space in black hole thermodynamics using the Iyer-Wald formalism. To proceed, we introduce the operator $\tilde{\delta}$, which includes variations of Newton's constant $G$, to distinguish it from the ordinary variation $\delta$. By analyzing the Lagrangian through the variation operator $\tilde \delta \mathbf{L}(\phi, G)$, we can derive
\begin{align}
	\tilde{\delta} \mathbf{L}=& \frac{\partial \mathbf{L}}{\partial \phi}\delta \phi+\frac{\partial \mathbf{L}}{\partial G}\tilde\delta G+\frac{\partial \phi}{\partial G}\tilde\delta G
    =\frac{\partial \mathbf{L}}{\partial \phi}\tilde{\delta} \phi+ \frac{\partial \mathbf{L}}{\partial G}\tilde\delta G.
\end{align}
The variation depends on both the explicit dependence of $\mathbf{L}$ on $\phi$ and $G$, and the implicit dependence arising from $\phi$'s dependence on $G$. In the second equality, we utilize the relation $\tilde\delta \phi = \delta \phi + \frac{\partial \phi}{\partial G}\tilde\delta G$. This transformation simplifies the variation expression to
\begin{align}
	\begin{split}
		\tilde{\delta} \mathbf{L}
    =\mathbf{E}^{\phi} \tilde{\delta}  \phi+  d \boldsymbol{\Theta}[\tilde{\delta} \phi]+F_g\tilde{\delta}G\boldsymbol{\epsilon},
	\end{split}
\end{align}
where $F_g=(\frac{\partial \mathbf{L}}{\partial G})_{g_{\mu\nu}}$. Moreover, by rederiving Equation (\ref{intfor1}) using the Killing vector $\xi_H$ at the horizon and integrating over a hypersurface $\Sigma$, it yields
\begin{align}
	\begin{split}
		\int_{S_r}\left(\tilde{\delta} \mathbf{Q}_{\xi_H} -  {\xi_H}\cdot \boldsymbol{\Theta}[\tilde{\delta} \phi] \right) - \int_{S_H} \tilde{\delta} \mathbf{Q}_{\xi_H}=-\tilde{\delta}G \int_{\Sigma}F_g{\xi_H}\cdot\boldsymbol{\epsilon} .\label{SAdS}
	\end{split}
\end{align}
In what follows, we present specific calculation results for several typical black hole models.

\subsection{Schwarzschild-AdS black hole}
We start with the simplest model of a Schwarzschild-AdS (SAdS) black hole. The bulk action is given by
\begin{align}
	I=\frac{1}{16\pi G}\int_{\mathcal{M}}d^4x \sqrt{-g}(R-2\Lambda),
\end{align}
with the negative cosmological constant $\Lambda=-3/l^2$ related to the AdS radius $l$ and Newton's constant $G$. The metric of the SAdS black hole takes the form
\begin{align}
	d s^{2}&=-f(r) d t^{2}+\frac{1}{f(r)} d r^{2}+r^{2}(d\theta^{2}+\sin^2{\theta}d\phi^{2}),
\end{align}
with the metric function
\begin{align}
	f(r)=1-\frac{2GM}{r}+\frac{r^2}{l^2},
\end{align}
where $M$ is the mass of the black hole, and we denote $r_h$ solving from $f (r_h) = 0$ as the radius of the event horizon.

For the SAdS black hole, the time-like Killing vector at the horizon is $\xi_H = \frac{\partial}{\partial t}$. Upon the variation of the Lagrangian of Einstein gravity, the total derivative is obtained
\begin{align}
	\Theta^\mu (\delta g_{\mu \nu})\equiv \frac{1}{16 \pi G} (g^{\mu \alpha} \nabla^\nu \delta g_{\alpha \nu} - g^{ \alpha \beta} \nabla^{\mu} \delta g_{ \alpha \beta} ), \label{einsteintheta}
\end{align}
and Noether charge is defined as \cite{Iyer:1994ys}
\begin{align}
	Q_\xi^{\mu\nu} \equiv -\frac{1}{16\pi G}  \left (  \nabla^\mu \xi^\nu -\nabla^{\nu}\xi^\mu  \right ), \label{einq}
\end{align}
here Newton's constant $G$ should be regarded as variable with respect to the symbol $\tilde{\delta}$. Substituting Eqs.\eqref{einsteintheta} and \eqref{einq} into the relation \eqref{SAdS} , it results in
\begin{align}
	\tilde{\delta} M+\frac{M}{2G}\tilde{\delta}G-\cancel{\frac{r^3}{2 G^2 l^2}\tilde{\delta} G} -T\tilde{\delta} S=-\cancel{\frac{r^3}{2 G^2 l^2}\tilde{\delta} G}+\frac{r_h^3}{2 G^2 l^2}\tilde{\delta} G.
    \label{delt}
\end{align}
To further analyze the thermodynamic relations of the black hole system, we invoke the holographic dictionary\cite{Myers:2010tj}, which establishes the relation $C \propto l^2/G$  for the central charge of the boundary field theory. In the restricted phase space of black hole thermodynamics, the quantity $C$ is treated as a new thermodynamic variable.
Then, the Equation (\ref{delt}) turns to
\begin{align} \label{delta M}
\widetilde{\delta}M=T\widetilde{\delta}S+(\frac{r_h}{4 l^2}-\frac{r_h^3}{4 l^4})\widetilde{\delta}C=T\widetilde{\delta}S+\mu_g\widetilde{\delta}C,
\end{align}
where $\mu_g$ is defined as the chemical potential conjugate to the central charge $C$.     

On the other hand, the chemical potential $\mu_g$ can be obtained from the Euclidean action of the bulk spacetime via $TI_E=\mu_g C$. 
The explicit form of the Euclidean action on the shell of the 4-dimensional SAdS black hole spacetime is given in \cite{york1986black}
\begin{align}
		T I_{E} =\lim _{r \rightarrow \infty}\left(I_{B H}-I_{A d S}\right)\nonumber=\frac{r_h}{4 G}-\frac{r_h^3}{4 l^2 G},
\end{align}
where the action $I_{BH}$ consists of contributions from the Euclidean bulk, the surface term, and the counterterms, and the subtraction action $I_{AdS}$ is introduced to remove the divergent terms at infinity.
It is easy to verify the formula for the chemical potential $\mu_g$
\begin{align}
		\mu_g=\frac{r_h}{4 l^2}-\frac{r_h^3}{4 l^4}.
\end{align}

Additionally, from Eq.\eqref{intfor2} we can obtain the homogeneous Euler relation corresponding to traditional thermodynamics
\begin{align}
	M=TS+\mu_g C.
\end{align}
Thus, by appropriately extending the Iyer-Wald formalism, we obtain the first law and the Euler relation
of thermodynamics for SAdS black holes within the restricted phase space formalism. 

\subsection{Reissner-Nordström-AdS black hole}
The action of 4-dimensional RN-AdS black holes is expressed as
\begin{align}
	I=\frac{1}{16\pi G}\int_{\mathcal{M}}d^4x \sqrt{-g}(R-2\Lambda)-\frac{1}{16 \pi}\int_{\mathcal{M}}d^4x (F_{\mu\nu}F^{\mu\nu}),
\end{align}
where $F_{\mu \nu}= \partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the Maxwell field tensor. The RN-AdS black hole solution is
\begin{align}
	d s^{2}&=-f(r) d t^{2}+\frac{1}{f(r)} d r^{2}+r^{2}(d\theta^{2}+\sin^2{\theta}d\phi^{2}),\\
	A_{\mu}&=(\Phi(r),0,0,0), 
\end{align}
the metric function and the charge potential are
\begin{align}
	f(r)=1-\frac{2GM}{r}+\frac{G Q^2}{r^2}+\frac{r^2}{l^2},~~~~~\Phi(r)=\frac{Q}{r}.
\end{align}
From the variation $\delta ( \sqrt{-g} L )$, one can get \cite{Hawking:1995ap}
\begin{align}
	\Theta^\mu(\delta g_{\alpha\beta}, \delta A_\nu)&\equiv \frac{1}{16\pi G} \left( g^{\mu\alpha} \nabla^\beta \delta g_{\alpha\beta}- g^{\alpha\beta} \nabla^\mu \delta g_{\alpha\beta} \right)- \frac{1}{4\pi G} F^{\mu\nu} \delta A_\nu,\\
	Q_\xi^{\mu\nu} &\equiv - \frac{1}{16\pi G} \left( \nabla^\mu \xi^\nu - \nabla^\nu \xi^\mu \right)- \frac{1}{4\pi G} \left( \xi^\rho A_\rho \right) F^{\mu\nu}.	
\end{align}
Substituting the above results into the Equation \eqref{SAdS} and conducting further simplifications, we arrive at the following expression
\begin{align}
	\widetilde{\delta}M=T\widetilde{\delta}S+\hat{\Phi}\widetilde{\delta}\hat{Q}+\left(\frac{r_h}{4 l^2}-\frac{r_h^3}{4 l^4}-\frac{G Q^2}{4 r_h l^2}\right)\widetilde{\delta}C.
\end{align}
Where $\hat{\Phi}=\frac{\Phi(r_h) \sqrt{G}}{l}, \hat{Q}=\frac{Q l}{\sqrt{G}}$ are the rescaled potential and charge respectively due to different conventions in the action\cite{Cong:2021fnf}. The explicit expression for the Euclidean action on the shell of the 4-dimensional RN-AdS black hole spacetime is calculated as follows\cite{Chamblin:1999tk}
\begin{equation}
	\mu_g C=TI_E=\frac{r_h}{4 G}-\frac{r_h^3}{4 l^2 G}-\frac{Q^2}{4 r_h},
\end{equation}
with the definition $C=l^2/G$, we promptly obtain
\begin{align}
	\widetilde{\delta}M=T\widetilde{\delta}S+\hat{\Phi}\widetilde{\delta}\hat{Q}+\mu_g\widetilde{\delta}C.
\end{align}
Subsequently, from Eq.\eqref{intfor2} the homogeneous Euler relation is obtained
\begin{align}
	M=TS+\hat{\Phi} \hat{Q}+\mu_g C.
\end{align}
Therefore, when the charge $Q$ exists, we use a geometric approach to construct a homogeneous black hole thermodynamic framework, which is consistent with the results in \cite{Zeyuan:2021uol}.
\subsection{Kerr-AdS black hole}
The Kerr-AdS black hole is a stationary solution of Einstein gravity and has a distinct geometric structure, making it worthy of in-depth study. Its metric can be expressed as follows
\begin{align*}
	d s^2= & -\frac{\Delta_r}{\rho^2}\left(d t-\frac{a \sin ^2 \theta}{\Xi} d \phi\right)^2+\frac{\rho^2}{\Delta_r} d r^2+\frac{\rho^2}{\Delta_\theta} d \theta^2+\frac{\Delta_\theta \sin ^2 \theta}{\rho^2}\left(a d t-\frac{r^2+a^2}{\Xi} d \phi\right)^2,
\end{align*}
each of these items are defined as
\begin{align*}
		\rho^2&=r^2+a^2 \cos ^2 \theta,~~~~~\Delta_r=\left(r^2+a^2\right)\left(1+\frac{r^2}{l^2}\right)-2 G m r, \\
\Xi&=1-\frac{a^2}{l^2}, ~~~~~~~~~~~~~~\Delta_\theta=1-\frac{a^2}{l^2} \cos ^2 \theta,
\end{align*}
where $a$ is the rotation parameter of the black hole. The mass $M$ and angular momentum $J$ are related to the parameters $m, a$ via the following expressions \cite{Gibbons:2004ai}
\begin{align}
M=\frac{m}{\Xi^2},~~~~~J=	\frac{a m}{\Xi^2},  \label{mj}
\end{align}
in which $m$ can be solved from the equation $\Delta_r(r_h) = 0$. In Kerr-AdS spacetime, the horizon Killing vector decomposes into two parts, i.e., $\xi_H=\xi_t+\Omega\, \xi_\phi$, where $\xi_t \equiv \frac{\partial}{\partial t}+\Omega_{\infty} \frac{\partial}{\partial \phi}$ is the modified time-like Killing vector that accounts for the angular velocity at infinity, $\xi_\phi\equiv \frac{\partial}{\partial \phi}$ is the rotational Killing vector, and  $\Omega\equiv \Omega_H-\Omega_\infty$ represents the angular velocity measured relative to infinity.

This decomposition of the horizon Killing vector makes it straightforward to see that
\begin{align}
	\textbf{Q}_{\xi_H} = \textbf{Q}_{\xi_t}+\textbf{Q}_{\xi_\phi}.
\end{align}
With the definition of the Noether charge (\ref{einq}), substitute it into Equation \eqref{SAdS}, which yields
\begin{align}
 \frac{l^4}{(a^2-l^2)^2}\widetilde{\delta}m-\frac{4al^4m}{(a^2-l^2)^3}\widetilde{\delta}a
 -\frac{l^2 m}{2G(a^2-l^2)}\widetilde{\delta} G
 -\Omega \widetilde{\delta} J 
 -T \widetilde{\delta} S
 = \frac{r_h(a^2+r_h^2)}{2G^2(l^2-a^2)}\widetilde{\delta} G.
\end{align}
By consideration the relations Eqs.(\ref{mj}), we find that
\begin{align}
 \frac{l^4}{(a^2-l^2)^2}\widetilde{\delta}m-\frac{4al^4m}{(a^2-l^2)^3}\widetilde{\delta}a
  =\widetilde{\delta}M,
\end{align}
which simplify the relation to
\begin{align}
 \widetilde{\delta}M&=T\widetilde{\delta}S+\Omega \widetilde{\delta}J+\frac{(l^2-r_h^2)(a^2+r_h^2)}{4l^2 r_h(a^2-l^2)}\widetilde{\delta}C.
\end{align}

Similarly, solving for the chemical potential $\mu_g$ from the on-shell action \cite{Gibbons:2004ai}, we find that 
$$
    \mu_g=\frac{(l^2-r_h^2)(a^2+r_h^2)}{4l^2 r_h(a^2-l^2)},
$$ 
which constructs the relation
\begin{align}
	\widetilde{\delta}M&=T\widetilde{\delta}S+\Omega \widetilde{\delta}J+\mu_g\widetilde{\delta}C,\\
	M&=TS+\Omega J+\mu_g C.
\end{align}
In summary, although the Kerr-AdS black hole has a more complex structure, we have obtained precise expressions for the relevant thermodynamic quantities through the extended Iyer-Wald formalism, as well as the Euler relation consistent with the first law. This agrees with the results presented in \cite{Gao:2021xtt}.

\section{Conclusion}\label{IV}
In this work, we have extended the Iyer-Wald formalism and constructed a self-consistent and universal framework for homogeneous black hole thermodynamics in any diffeomorphism-invariant gravitational theory. This framework derives explicit expressions for all thermodynamic variables directly from geometric principles (Noether currents and charges arising from the Lagrangian density), rather than assuming thermodynamic laws a priori or directly deriving them from thermodynamic relations. In the restricted phase space where the cosmological constant $l$ is fixed and Newton’s constant $G$ varies, the central charge $C$ and its conjugate chemical potential $\mu$ as geometrically defined thermodynamic variables are introduced. This extension enables us to rigorously establish the first law of thermodynamics and the Euler homogeneous relation, thereby ensuring the extensivity essential for alignment with classical thermodynamics.

We explicitly demonstrate the consistency of this framework through applications to three canonical black hole solutions, i.e., Schwarzschild-AdS, Reissner-Nordström-AdS, and Kerr-AdS. In each case, we derive explicit thermodynamic equations and variable expressions. Crucially, we independently compute the chemical potential using the gravitational Euclidean action on the shell and find perfect agreement with the geometrically derived values. 

The profound universality of the Iyer-Wald formalism shows its potential as a foundational tool for constructing black hole thermodynamics. Furthermore, it is worth noting that the precise thermodynamic roles of a-charges and c-charges remain incompletely understood, where these charges decouple \cite{Li:2018drw}, the extended Iyer-Wald formalism may offer a powerful geometric path to clarify these roles. We anticipate broad applications ultimately revealing richer black hole thermodynamic phenomena that bridge quantum gravity, information theory, and the emergence of spacetime. 
\section*{Acknowledgement}
\vspace{-0.7em}
This work is supported by the National Natural Science Foundation of China (Grant Nos.12275216, 12105222, 12247103).
\end{spacing}

\end{document}